# Selective Prompt Anchoring for Code Generation

### v4

@@ 5.4 Ablation Study of Anchored Text Selection
To investigate the impact of anchored text selection in code generation tasks, we conduct an ablation study by masking different components in code generation prompts. We decompose the code generation prompt into three components: (1) natural language description, (2) source code, and (3) test cases. We create 4 automated anchored text selection methods by ablating the source code and test cases.777We do not ablate the natural language description since it explicitly represents the user intent. To explore whether anchoring on a few more informative tokens can enhance performance, we further create a condition using important tokens labeled by human programmers as the anchored text. We leverage the dataset from Kou et al. [23], where multiple programmers manually identified critical tokens that models should attend to when solving HumanEval and MBPP programming tasks. Following our previous experimental setup, we tune Spa for each condition and benchmark, calculating the average Pass@ improvement across all 6 experimental LLMs.
@@
| Anchored Text | He | He+ | Mb | Mb+ | Bcb | ||
3.66 | 3.04 | 2.58 | 2.11 | n/a | ||
| Natural Language | 5.48 | 5.08 | 4.26 | 3.22 | 1.57 | ||
| + Code | 4.87 | 4.65 | 3.75 | 2.81 | 1.42 | ||
| + Test Cases | 5.11 | 4.89 | 4.05 | 3.11 | 1.50 | ||
| + Code & Test Cases | 4.76 | 4.57 | 3.98 | 2.81 | 1.44 |
Interestingly, we find that anchoring using human-labeled important tokens achieves only 3.66% on HumanEval, which is consistently less effective than all the automated experimental anchoring methods. It suggests that narrowing the anchored text to more informative tokens does not necessarily improve performance. We think there are two plausible reasons. First, since LLMs need to attend to different context tokens at each decoding step, providing a narrow set of anchored tokens may have a negative impact and distract the LLM in certain decoding steps. Second, previous studies such as Xiao et al. [57] show that even though some tokens, such as separators and empty space, may not be semantically meaningful or informative, they provide important signals for LLMs to generate the right content (e.g., following the grammar rules). Thus, over-attending the informative tokens but not the special tokens in the task description may disrupt the regular generation process. Nevertheless, we think this is a challenging but interesting future direction to investigate. We believe our findings will open up new research opportunities for the community on this topic.
5.5 Analysis of the Impact of Prompt Length
@@
To further validate that Spa effectively addresses the attention dilution issue, we analyze both the original model performance and Spa’s effectiveness across different prompt lengths, as shown in Table 5. We divide the HumanEval dataset into three equal-sized subsets (Short, Medium, and Long) based on the 33rd and 66th percentiles of prompt lengths. We find original code LLMs consistently achieve better performance on shorter prompts compared to longer ones. The average Pass@1 is 74.3% for Short prompts, 49.2% for Medium prompts, and only 30.8% for Long prompts. This result echoes our empirical finding on attention dilution, as longer prompts can lead to more severe attention dilution issues.
@@ 5.6 Evaluation on Other Generative Tasks
While we focus on code generation in this work, we are interested in whether Spa can be applied to other generative tasks, Thus, we evaluate Spa on other generative benchmarks, including TruthfulQA [28], GSM8K [8], MMLU [19], and BoolQ [7].
@@
As shown in Table 6, Spa enhances the base model’s performance in all benchmarks except MMLU. While the performance improvement is not as significant as in code generation, we hypothesize that this is because different tasks have unique input and output patterns, leading to varying degrees of attention dilution. For MMLU, the model only needs to generate a choice to answer the multiple-choice question. The output length is significantly shorter than in code generation and considerably less than in user prompts. Consequently, the model’s attention is hardly diluted by self-generated tokens, making Spa not helpful. In contrast, TruthfulQA requires the model to generate a text analysis, provide reasoning, and answer the question. Therefore, Spa is more beneficial in addressing attention dilution and correcting the remaining 27% errors for “Truth.” (Truthfulness) and 41% errors for “Info.” (Informativeness). Nevertheless, it remains an interesting future work to investigate how to further improve Spa for a broader range of generative tasks beyond code generation. More details are reported in Appendix I.
@@ 6 Related Work
Code Generation. To enhance the performance of LLMs on coding tasks, significant efforts have been dedicated to curating high-quality training data [26, 17, 56] and designing domain-specific training objectives [38, 2]. Additionally, techniques such as instruction tuning [55], reinforcement learning with human feedback [39], and repository-level context modeling [59] have been explored to improve alignment, reasoning, and context understanding abilities in code generation. However, these approaches demand substantial computational resources. To address this challenge, researchers have increasingly focused on prompting methods [4, 61, 47, 24], which is training-free and aims to optimize the prompt with more information. For example, Self-Debugging [4] and Self-Edit [61] enable LLMs to debug code based on error messages and execution results. Self-Planning [21] allows LLMs to decompose tasks into subtasks and implement solutions step-by-step. ReAct [58] prompts an LLM to generate reasoning traces and action plans in an interleaved manner. Compared with these methods, Spa takes an orthogonal approach that amplifies the influence of the user prompt to mitigate the attention dilution issue.
@@
Attention Steering. TOAST [44] tunes a feature selection module to redirect attention to task-relevant features. PASTA [62] performs model profiling to identify beneficial attention headers and recalculate attention distributions across transformer layers. However, these attention-steering approaches usually require extensive model adaptations and complex setup procedures. In contrast, Spa provides a model-agnostic solution that requires tuning only a single hyperparameter. Furthermore, unlike PASTA, which requires manual user input to guide attention, Spa automatically enhances code generation by amplifying the influence of the user prompt without humans in the loop.
@@
Logit Arithmetic. There has been a growing body of research on performing arithmetic transformations on logits to enhance text generation, such as contrasting logits from multiple LMs [30, 29, 11, 63] and contrasting logits from different layers of a model [6, 16]. Unlike these methods, Spa contrasts logits from the same model by perturbing the input through masking, rather than providing additional context [40, 45, 35] or changing to a completely new prompt [43]. Furthermore, we provide a mathematical approximation of semantic scaling over arbitrary groups of embeddings. Spa is specifically designed to address the attention dilution issue in LLMs during code generation—a phenomenon first observed in our work. By contrast, none of the existing works explored code generation or model attention. They primarily focus on reducing hallucinations [45, 43, 25], enhancing coherence [35], factuality [6], and controllable text generation [29, 40, 63].
@@ 7 Limitations & Future Work
While Spa achieves a consistent improvement on LLMs with different sizes and types, we do not observe a monotonic relationship between model attributes and the improvement. Furthermore, there is no obvious correlation between the original model performance and the improvement. It is an interesting future direction to investigate how different model attributes affect the improvement achieved by Spa. In this work, we pre-define the method for selecting anchored tokens and use a fixed anchoring strength when generating code. We consider this approach a baseline. Future work could explore dynamically determining both the anchored text and the anchoring strength at different steps. Furthermore, the underlying principle of Spa is not restricted to transformer-based LLMs and could be adapted for model architectures such as RNN.
@@ 8 Conclusion
We present Spa, a model-agnostic approach that enhances LLM code generation ability by addressing the attention dilution phenomenon, which we are the first to identify and characterize. Our empirical study indicates that LLMs tend to overlook the prompt as generating more new tokens. To mitigate this issue, Spa introduces a novel, training-free approach for controlling the contextual influence of selected prompt tokens. We mathematically prove that this approach can effectively steer model attention toward specified text segments. By selectively amplifying prompt influence through Spa, LLMs overcome the attention dilution issue and achieve significant, consistent improvements in code generation performance across multiple models and benchmarks.
A Gradient-based Attention Analysis A
B Attention Calculations in Empirical Study B
@@ G Hypothetical explanation for Attention dilution and SPA’s effectiveness G
H Setup of Large Language Models H
@@ I Evaluating Spa on Other Generative Tasks I
J Tuning of Anchoring Strength I
J.1 Analysis of Anchoring Strength J.2
J.2 Tuning Stability J.2
J.3 Cross-dataset & Cross-Model Tuning Evaluation J.3
J.4 Comparision between Tuned Spa and Optimal Spa J.4
@@
L Performance Analysis by Different Generation Lengths M
Appendix A Gradient-based Attention Analysis
To validate the generalizability of attention dilution, we employ a gradient-based attention calculation method [42, 46]. Unlike self-attention analysis, this method treats the entire LLM as a differentiable function, allowing us to measure each input token’s contribution to the output through gradient computation. We include calculation details in Appendix B.
@@ Appendix A Gradient-based Attention Analysis
Appendix B Attention Calculations in Empirical Study
Self-attention. Most LLMs are based on the decoder of transformer [50] which has multiple self-attention layers. Roughly speaking, given an LLM and an input sequence of tokens where represents the th token. The transformer calculates relevance scores between every pair of tokens. The self-attention score for a token in the sequence can be roughly formulated as:
where the relevance function approximates the computation among in transformers [50]. However, different layers have different attention distributions. According to a study [52], deeper self-attention layers can better capture long-distance dependencies and program structure, so we calculate the attention by aggregating attention from multiple heads at the last layer. Nevertheless, this still excludes the influence from the last forward layer.
Gradient-based Attention. Compared to using self-attention layers in transformers, the gradient-based method can be generalized to different model architectures and consider the entire model as a whole. It computes the model’s attention by calculating the gradients relative to the input. Intuitively, a token that induces a larger gradient is considered more influential, suggesting that the model pays greater attention to it. Formally, the attention over the token is calculated by
Attention Percentage to the Prompt. Based on these two methods, we analyze how the attention of LLMs to the initial prompt shifts. Formally, given the prompt and the following generated tokens , we calculate the percentage of attention over the initial prompt
Given attention analysis requires open sourcing, we select five SOTA code LLMs with various sizes. We run the experiments on HumanEval [3], one of the most popular benchmarks for evaluating code generation models. We run five LLMs [36, 41, 17] on all 164 Humaneval tasks. Figure 4 shows the gradient-based attention shift when generating the first 400 tokens. The value gradually becomes noisy due to the lack of generated sequence with enough length.
The results demonstrate that there indeed exists such attention dilution issue. Due to the autoregressive nature, LLMs’ attention to the initial prompt is gradually diluted as generating more code. LLMs tend to attend to code generated by itself. Our finding is supported by another study [5] which investigates the self-attention dilution of transformers in a more general scenario.
@@ Appendix C Extended Discussion of Approximation in Spa
In Section J.1, our experiments reveal that ’s impact on code generation performance follows an unimodal pattern—initially increasing, then decreasing. Due to its distribution simplicity, we argue that while a higher-order approximation may yield a more reasonable performance distribution across different values, it does not significantly affect the process of locating the optimal anchoring strength. Therefore, beyond its computational efficiency, the first-order approximation in Spa is adequate for calculating semantically accurate augmented logits.

### v5

@@ 5.4 Ablation Study of Anchored Text Selection
To investigate the impact of anchored text selection in code generation tasks, we conduct an ablation study by masking different components in code generation prompts. We decompose the code generation prompt into three components: (1) natural language description, (2) source code, and (3) test cases. We create 4 automated anchored text selection methods by ablating the source code and test cases.888We do not ablate the natural language description since it explicitly represents the user intent. To explore whether anchoring on a few more informative tokens can enhance performance, we further create a condition using important tokens labeled by human programmers as the anchored text. We leverage the dataset from Kou et al. (2024), where multiple programmers manually identified critical tokens that models should attend to when solving HumanEval and MBPP programming tasks. Following our previous experimental setup, we tune Spa for each condition and benchmark, calculating the average Pass@ improvement across all 6 experimental LLMs.
| Anchored Text | He | He+ | Mb | Mb+ | Bcb | Lcb | ||
3.66 | 3.04 | 2.58 | 2.11 | n/a | n/a | ||
| Natural Language | 5.48 | 5.08 | 4.26 | 3.22 | 1.57 | 1.90 | ||
| + Code | 4.87 | 4.65 | 3.75 | 2.81 | 1.42 | 1.73 | ||
| + Test Cases | 5.11 | 4.89 | 4.05 | 3.11 | 1.50 | 1.82 | ||
| + Code & Test Cases | 4.76 | 4.57 | 3.98 | 2.81 | 1.44 | 1.73 |
@@
Interestingly, we find that anchoring using human-labeled important tokens achieves only 3.66% on HumanEval, which is consistently less effective than all the automated experimental anchoring methods. It suggests that narrowing the anchored text to more informative tokens does not necessarily improve performance. We think there are two plausible reasons. First, since LLMs need to attend to different context tokens at each decoding step, providing a narrow set of anchored tokens may have a negative impact and distract the LLM in certain decoding steps. Second, previous studies such as Xiao et al. (2024) show that even though some tokens, such as separators and empty space, may not be semantically meaningful or informative, they provide important signals for LLMs to generate the right content (e.g., following the grammar rules). Thus, over-attending to the informative tokens but not the special tokens in the task description may disrupt the regular generation process. Nevertheless, we think this is a challenging but interesting future direction to investigate. We believe our findings will open up new research opportunities for the community on this topic.
5.5 Impact of Prompt Length
To further validate that Spa effectively addresses the attention dilution issue, we analyze both the original model performance and Spa’s effectiveness across different prompt lengths, as shown in Table 6. We divide the HumanEval dataset into three equal-sized subsets (Short, Medium, and Long) based on the 33rd and 66th percentiles of prompt lengths. We find that the original code LLMs consistently achieve better performance on shorter prompts compared to longer ones. The average Pass@1 is 74.3% for Short prompts, 49.2% for Medium prompts, and only 30.8% for Long prompts. This result echoes our empirical finding on attention dilution, as longer prompts can lead to more severe attention dilution issues.
@@ 5.5 Impact of Prompt Length
5.6 Impact of Anchoring Strength
In this section, we study how the anchoring strength of Spa affects performance. To ensure fair comparisons, Spa is activated for all tasks in these experiments, unlike during inference when it is only activated for failed tasks. We also tune Spa on all tasks to improve generalizability. As shown in Figure 3, represents the original model.
We find that different anchoring strengths have different effects: if is too low, improvements are low; if too high, the LLM starts becoming biased and performance declines. Overall, as increases, performance first improves, reaches an optimum, and then drops with further increases.
We find is both tunable and transferable. While the optimal value varies slightly across models and benchmarks, it is generally model-dependent. We find serves as a universal anchoring strength, consistently improving performance in all cases. Additional experiments and discussions are provided in Appendix J.
@@ 5.7 Evaluation on Other Generative Tasks
While we focus on code generation in this work, we are interested in whether Spa can be applied to other generative tasks, Thus, we evaluate Spa on other generative benchmarks, including TruthfulQA (Lin et al., 2022), GSM8K (Cobbe et al., 2021), MMLU (Hendrycks et al., 2021b), and BoolQ (Clark et al., 2019). As shown in Table 7, Spa enhances the base model’s performance in all benchmarks except MMLU. While the performance improvement is not as significant as in code generation, we hypothesize that this is because different tasks have unique input and output patterns, leading to varying degrees of attention dilution. For MMLU, the model only needs to generate a choice to answer the multiple-choice question. The output length is significantly shorter than in code generation and considerably less than in user prompts. Consequently, the model’s attention is hardly diluted by self-generated tokens, making Spa not helpful. In contrast, TruthfulQA requires the model to generate a text analysis, provide reasoning, and answer the question. Therefore, Spa is more beneficial in addressing attention dilution and correcting the remaining 27% errors for “Truth.” (Truthfulness) and 41% errors for “Info.” (Informativeness). Nevertheless, it remains an interesting future work to investigate how to further improve Spa for a broader range of generative tasks beyond code generation. More details are reported in Appendix I.
@@ 6 Related Work
Code Generation. To enhance the performance of LLMs on coding tasks, significant efforts have been dedicated to curating high-quality training data (Li et al., 2023; Guo et al., 2024; Tian et al., 2025; Wei et al., 2023) and designing domain-specific training objectives (Niu et al., 2022; Chakraborty et al., 2022). Furthermore, techniques such as instruction tuning (Wei et al., 2022), reinforcement learning with human feedback (Ouyang et al., 2022), and repository-level context modeling (Zhang et al., 2023a) have been explored to improve alignment, reasoning, and context understanding abilities in code generation. However, these approaches often require significant fine-tuning effort. In the meantime, a line of work has focused on developing interactive methods (Di & Zhang, 2025; Tian et al., 2023, 2024) that incorporate real-time human feedback to guide and refine the model’s output. Despite their effectiveness, such human-in-the-loop approaches remain limited due to the requirement of user intervention. To overcome these challenges, recent research has increasingly explored automated and training-free prompting methods (Chen et al., 2024; Zhang et al., 2023b; Suzgun et al., 2022; Le et al., 2024), which aim to automatically optimize prompts by integrating additional reasoning or contextual information. For example, Self-Debugging (Chen et al., 2024) enable LLMs to debug code based on error messages and execution results. Self-Planning (Jiang et al., 2024) allows LLMs to decompose tasks into subtasks and implement solutions step-by-step. ReAct (Yao et al., 2023) prompts an LLM to generate reasoning traces and action plans in an interleaved manner. Compared with these methods, Spa takes an orthogonal approach that amplifies the influence of the user prompt to mitigate the attention dilution issue.
@@
Attention Steering. TOAST (Shi et al., 2023a) tunes a feature selection module to redirect attention to task-relevant features. PASTA (Zhang et al., 2024) performs model profiling to identify beneficial attention headers and recalculate attention distributions across transformer layers. However, these attention-steering approaches usually require extensive model adaptations and complex setup procedures. In contrast, Spa provides a model-agnostic solution that mathematically simulates attention steering via logit arithmetic.
@@
Logit Arithmetic. There has been a growing body of research on performing arithmetic transformations on logits to enhance text generation, such as contrasting logits from multiple LMs (Liu et al., 2024a, 2021; Dou et al., 2019; Zhao et al., 2024) and contrasting logits from different layers of a model (Chuang et al., 2024; Gera et al., 2023). Unlike these methods, Spa contrasts logits from the same model by perturbing the input through masking, rather than providing additional context (Pei et al., 2023; Shi et al., 2023b; Malkin et al., 2022) or changing to a completely new prompt (Sennrich et al., 2024). Furthermore, we provide a mathematical approximation of semantic scaling over arbitrary groups of embeddings. Spa is specifically designed to address the attention dilution issue in LLMs during code generation—a phenomenon first observed in our work. By contrast, none of the existing works explored code generation or model attention. They primarily focus on enhancing coherence (Malkin et al., 2022), factuality (Shi et al., 2023b; Chuang et al., 2024; Sennrich et al., 2024; Leng et al., 2023), and controllability (Liu et al., 2021; Pei et al., 2023; Zhao et al., 2024).
@@ 7 Limitations & Future Work
In this work, we pre-define the method for selecting anchored tokens and use a fixed anchoring strength when generating code. We consider this approach a baseline. Future work could explore how to dynamically select the anchored text and anchoring strength. For the selection of anchored text, one idea is to use LLMs to dynamically identify relevant words or phrases corresponding to the current generation step. For tasks where NL may not be important compared to code (e.g., code translation), we can use static code analysis to identify important code elements (e.g., function calls and variable names heavily used in the code). For the selection of anchoring strength, one idea is to develop a method to calculate the relevance of words and phrases to each generation step. Based on the relevance scores, the system can assign higher values to more relevant contexts while assigning lower values to less relevant ones.
In our current experiments, we utilize existing test cases in the benchmark to determine whether to activate Spa. Although this is a common practice in code generation (Chen et al., 2024; Zhang et al., 2023b), this may not be applicable to programming tasks where test cases are not available. This can be potentially addressed by prompting LLMs to generate test cases. Furthermore, we believe Spa will be beneficial in a self-improving pipeline for fairly complex tasks. By analyzing the errors in the initially generated code, LLMs can be prompted to identify which instructions or requirements were not followed in the prompt. Then Spa can be used to amplify the influence of these ignored instructions.
@@ 8 Conclusion
We present Spa, a model-agnostic approach to enhancing LLM code generation. Our study first identifies the attention dilution phenomenon, where code LLMs increasingly overlook the prompt as generation progresses. To address this, Spa introduces a training-free, mathematically proven mechanism for controlling the influence of selected prompt tokens. Experiments demonstrate that aligning model attention to user prompts using Spa significantly and consistently enhances the code generation performance of base LLMs.
Acknowledgements
We thank all the anonymous reviewers for their valuable and detailed feedback, which has significantly improved the quality of this paper. This work was supported in part by the National Science Foundation (NSF CAREER Grant 2340408).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
A Performance Analysis by Different Generation Lengths A
B Calculation Details and Extended Discussion of LLM Attention B
B.1 Calculation of Self-attention B.1
B.2 Calculation of Gradient-based Attention B.2
B.3 Attention Misalignment B.3
B.4 Attention to the User Prompt Ratio B.4
@@ E Computational Cost of Spa E
E.1 Latency E.1
E.2 Memory E.2
E.3 Tradeoff Between Latency and Memory Usage via Parallelization E.3
@@ G Hypothetical explanation for Attention dilution and SPA’s effectiveness G
H Implementation and Deployment H
H.1 Imlementation of Spa H.1
H.2 Model Deployment H.2
H.3 Prompt Design H.3
@@ I Evaluating Spa on Other Generative Tasks I
J Tuning of Anchoring Strength J
J.1 Tuning Stability J.1
J.2 Cross-dataset & Cross-Model Tuning Evaluation J.2
J.3 Comparision between Tuned Spa and Optimal Spa J.3
@@
Appendix A Performance Analysis by Different Generation Lengths
To evaluate how code length affects model performance, we analyze the length distribution of correctly and incorrectly generated code. We use the same set of experimental LLMs (Nijkamp et al., 2023; Guo et al., 2024; Rozière et al., 2024) as in Section 2 and experiment on both HumanEval (Chen et al., 2021) and LiveCodeBench (Jain et al., 2024).
Particularly, we evaluate code length at three granularities: characters, words (tokens), and lines. Figure 4 demonstrates that incorrectly generated code is significantly longer than correctly generated code across different granularity levels. To mitigate the impact of the correlation between task difficulty and generation length, we conduct further analysis at different task difficulty levels in LiveCodeBench. Table 1 shows that the average length of incorrectly generated code is consistently longer than that of correctly generated code, across all levels.
These results echo our findings about attention dilution—as the generated sequence grows longer, the model’s attention to the user prompt diminishes, leading to more errors.
Appendix B Calculation Details and Extended Discussion of LLM Attention
B.1 Calculation of Self-attention
Most LLMs are based on the decoder of transformer (Vaswani et al., 2017), which has multiple self-attention layers. Roughly speaking, given an LLM and an input sequence of tokens where represents the th token. The transformer calculates relevance scores between every pair of tokens. The self-attention score for a token in the sequence can be roughly formulated as:
where the relevance function approximates the computation among in transformers (Vaswani et al., 2017). However, different layers have different attention distributions. According to a study (Wan et al., 2022b), deeper self-attention layers can better capture long-distance dependencies and program structure, so we calculate the attention by aggregating attention from multiple heads at the last layer. Nevertheless, this still excludes the influence from the last forward layer.
B.2 Calculation of Gradient-based Attention
To validate the generalizability of attention dilution, we employ a gradient-based attention calculation method (Selvaraju et al., 2016; Shrikumar et al., 2017). Compared to using self-attention layers in transformers, the gradient-based method can be generalized to different model architectures by treating the entire model as a whole differentiable function. It computes the model’s attention by calculating the gradients relative to each input token. Intuitively, a token that induces a larger gradient is considered more influential, suggesting that the model pays greater attention to it. Formally, the attention over the token is calculated by
@@ B.2 Calculation of Gradient-based Attention
B.3 Attention Misalignment
Despite the success of the attention mechanism in LLMs, prior works found that language models often exhibit simple attention patterns (Raganato & Tiedemann, 2018; Voita et al., 2019). Furthermore, an empirical study (Kou et al., 2024) found that given a coding task, there often exists a misalignment between LLM attention and human attention. When generating code, LLMs often focus on parts in natural language descriptions that are different from what human programmers focus on. Another work (Ning et al., 2024) shows that when the model’s attention aligns more closely with human programmers’ attention, the model generates more accurate SQL queries. Inspired by these findings, we hypothesize that a root cause of inaccuracy in LLM-generated code stems from the suboptimal model attention.
B.4 Attention to the User Prompt Ratio
Based on these two methods to calculate LLMs’ attention, we analyze how the attention of LLMs to the initial prompt shifts. Formally, given the prompt and the following generated tokens , we calculate the attention to the user prompt ratio over the initial prompt
Given that attention analysis requires open sourcing, we select five SOTA code LLMs with various sizes. We run the experiments on HumanEval (Chen et al., 2021), one of the most popular benchmarks for evaluating code generation models. We run five LLMs (Nijkamp et al., 2023; Rozière et al., 2024; Guo et al., 2024) on all 164 Humaneval tasks. Figure 5 and Figure 2 show the gradient-based attention shift when generating the first 400 tokens. The value gradually becomes noisy due to the insufficient length of the generated sequence.
The results demonstrate that there indeed exists such an attention dilution issue. Due to the autoregressive nature, LLMs’ attention to the initial prompt is gradually diluted as they generate more code. LLMs tend to attend to code generated by themselves. Our finding is supported by another study (Chiang & Cholak, 2022) which investigates the self-attention dilution of transformers in a more general scenario.
@@ Appendix C Extended Discussion of Approximation in Spa
In Section 5.6, our experiments reveal that ’s impact on code generation performance follows an unimodal pattern—initially increasing, then decreasing. Due to its distribution simplicity, we argue that while a higher-order approximation may yield a more reasonable performance distribution across different values, it does not significantly affect the process of locating the optimal anchoring strength. Therefore, beyond its computational efficiency, the first-order approximation in Spa is adequate for calculating semantically accurate augmented logits.